# Finding Stable Price Zones in European Electricity Markets: Aiming to Square the Circle?

### v2

@@ Abstract
The European day-ahead electricity market is split into multiple bidding zones. Within these zones, a uniform energy price is computed for each hour. Large bidding zones have been under scrutiny. The fact that zonal clearing ignores the transmission capacities within zones and the increase in renewables lead to a growing number of interventions in the generation of energy sources and large redispatch costs. The European Union Agency for the Cooperation of Energy Regulators (ACER) proposed alternative bidding zone configurations that should be analyzed as part of the Bidding Zone Review. For Germany, four alternative configurations were suggested. Bidding zones shall be stable and based on long-term, structural congestion in the grid. We analyzed the proposed configurations considering different clustering algorithms and periods. We found that the configurations do not reduce the price standard deviations within zones considerably, while the average prices across zones are similar. Other configurations identified based on clustering prices lead to lower price standard deviations but are not geographically coherent. Importantly, different configurations emerge depending on clustering features, algorithm, and period considered. Given the substantial changes in energy supply and demand that can be expected in the future, defining stable configurations appears to be a moving target.
@@ 1 Introduction
The European electricity market is split into multiple bidding zones (BZs), aka. price zones. Within these zones, market participants can submit bids and offers, which are subsequently matched without considering the capacities of the transmission network. [^1] We will use the terms price zone and bidding zone interchangeably. BZs are generally defined depending on national borders, such that a country represents a single bidding zone (e.g., France and Poland). However, there are also countries that are divided into several BZs, such as Norway, Sweden and Italy, or BZs that consist of more than one country, like the German-Luxembourgish bidding zone.
@@
An important characteristic of BZs is that in the day-ahead market, they implement a uniform pricing mechanism, meaning that a single market-clearing electricity price applies to all participants belonging to a BZ. A uniform price in a BZ promotes transparency due to a simple single price signal in a large region, and it limits the market power that individual parties can exercise. However, a uniform electricity price also leads to significant negative ramifications. There are no local incentives to adapt to changes in the volatile supply of renewables, there are significant welfare losses due to the uniform price constraint, and large redispatch costs. The latter is because transmission constraints are ignored within a BZ when supply is matched with demand (Energywende, 2018), and the market-based allocation is not physically feasible. In addition, uniform prices do not set appropriate investment incentives.
@@
The German-Luxembourgish BZ is a case in point. Strong winds along the North Sea and Baltic Coast often contribute to elevated power production from wind farms in Northern Germany. The energy produced in the north must be transported to the main power consumption areas, which are located in the south and west of the country. However, the transmission lines connecting these regions often become congested and thus electricity cannot be delivered as determined in the day-ahead market. To stabilize the grid, redispatch by the Transmission System Operators (TSOs) is required. TSOs often need to throttle the generation of Northern conventional and wind power plants. At the same time, they instruct power plants in the south to increase energy production. Such redispatch actions incur substantial costs. In 2022, the redispatch costs in Germany were 4.2 billion Euros. [^2] https://www.bundesnetzagentur.de/
Furthermore, the lack of sufficient transmission lines in the German grid caused problems in neighboring countries. Poland and the Czech Republic have installed phase-shifters at their borders with Germany to reduce loop flows in their networks. [^3] https://www.cleanenergywire.org/news/southern-and-western-states-oppose-splitting-germanys-power-price-zone These loop flows lead to suboptimal cross-zonal trading opportunities and are a negative consequence of the large German bidding zone and its limited transmission grid.
@@
[…]
@@
In this paper, we leverage the data provided by ENTSO-E and study how stable and robust the configurations of price zones (or, clusterings) proposed by ACER are with respect to different time frames, different inputs (i.e., nodal prices with or without geographic coordinates), and different clustering algorithms. We define three criteria for analyzing clustering stability and robustness and conduct a thorough empirical evaluation of the suggested configurations using the data published by ENTSO-E. Following ACER’s methodology, we implement K-Means and Spectral Clustering and evaluate the clusterings computed with these methods. We want to understand if the clustering algorithms compute the same clusters when nodal prices of different time periods are considered. How different are clusters resulting from different clustering algorithms? Also, how similar are the clusters of a particular configuration with respect to prices, and how large are the price standard deviations within clusters?
@@
We find that the configurations proposed by ACER are not stable across time and algorithm used. Moreover, the configurations do not reduce the price standard deviations within zones considerably, and the average prices between zones are similar. If we recompute the clusters and take only prices into account [^6] This means that we do not include constraints in the clustering computation such that we obtain balanced clusters in terms of either the number of nodes or geographical coverage., we can reduce the price standard deviation within clusters more. However, these alternative clusters are not of equal size and are not geographically coherent. Importantly, neither the configurations proposed by ACER nor the clusters that we computed are stable across time. Depending on the time frame taken into account, different clusterings would result. Note that our analysis is performed based on the same target year 2025 as in the BZR. Given the significant changes in supply and demand and in the weather over time, we can expect substantial changes in the supply in future years. In particular, Germany has ambitious goals concerning the share of renewable energies and aims for 80% in 2030.777https://www.energiewechsel.de/KAENEF/Navigation/DE/Energiewechsel/Erneuerbare-Energien/erneuerbare-energien.html There is a danger that even if a good configuration is selected for 2025, it might be outdated soon after.
@@ 2 Related Literature
This paper complements existing literature computing bidding zone configurations. In particular, we draw on a unique data set with LMP prices that is provided by ENTSO-E. Bovo et al., (2019) identifies the main network indicators that should be considered for defining suitable bidding zone configurations. LMPs and power transfer distribution factors (PTDFs) represent such indicators. While some papers focus on identifying configurations by clustering PTDFs (e.g., (Kumar et al., 2004) and (Kłos et al., 2014)), most works use LMPs as clustering features (Stoft, 1997). Burstedde, (2012) computes LMPs for the scenario years 2015 and 2020 considering a simplified 72-node representation of the European grid, and evaluates the shapes and sizes of the zones obtained by clustering these LMPs using hierarchical clustering. Breuer et al., (2013) implement a genetic algorithm and apply it on a dataset consisting of 380 kV nodes considering a projected system for 2016 and 2018. Based on six scenarios, Felling & Weber, (2018) compute robust price zone configurations for Central Western Europe by employing hierarchical clustering based on the LMPs of approximately 2,200 nodes and assigning weights to nodes depending on their demand- and supply situation. These configurations are robust in the sense that LMPs from multiple scenarios are integrated into their computation. Considering a network model consisting of around 2,700 nodes, Brouhard et al., (2023) design nine zone delineations for Continental Europe by implementing K-Means and hierarchical clustering. The authors calculate LMPs for multiple, independent single-period optimization problems (i.e., optimal power flow with direct current approximation), and use geographic coordinates and these prices as clustering features. Sarfati et al., (2015) introduce five indicators to assess the impact of bidding zone configurations in zonally-priced electricity markets. The indicators include commercial exchange and welfare evolution, as well as price convergence and divergence, and were analyzed on the Nordic 32-bus example system.
@@
In terms of the German-Luxembourgish bidding zone, Zinke, (2023) identifies north-south splits using hierarchical clustering based on the LMPs associated to approximately 500 nodes that are computed within a linear market and grid model for four scenario years. The author considers uncertain factors such as short-term weather patterns and long-term system changes to evaluate the potential decrease in redispatch costs resulting from such splits. Moreover, Ambrosius et al., (2020) consider a scenario for the year 2035 and determine welfare-maximizing price zones and available transfer capacities for a simplified representation of the German market area consisting of 28 nodes. Such zones are identified based on a mixed-integer nonlinear model that incorporates a graph partitioning problem on the first level. The prior literature used different data sources for these analyses.
@@ 3.1 Clustering Algorithms
Following ACER’s approach to identify alternative price zone configurations, we consider two clustering algorithms in this paper. ACER used three clustering approaches (ACER, 2022b ), i.e., K-Means, Spectral Clustering with constrained K-Means and Spectral Clustering with CDIRC, but the implementation details are not publicly available. Therefore, it is impossible to reproduce the exact algorithmic configuration ACER implemented. Instead, we consider the same family of clustering methods to evaluate alternative price zone configurations. In what follows, we briefly describe the clustering algorithms we implement.
3.1.1 K-Means
@@
3.1.2 Spectral Clustering
@@ 3.2.2 Temporal Stability
While geographic coordinates are static across time, prices are time-sensitive. Therefore, we are interested to see if pricing zones are stable over time. To decide this, we compare two clusterings $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$ that are computed for two different time periods with the same algorithm and considering the same features. For this comparison, we use the Rand index, which measures the similarity between $\mathcal{C}_{1},\mathcal{C}_{2}$ considering all pairs of data points and counting the pairs that are assigned to the same or to different clusters in both $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$:
@@ 4 Data
The data we use in our experiments was published as part of the locational pricing study conducted by ENTSO-E [^10] The datasets can be accessed at https://www.entsoe.eu/network_codes/bzr/ (January 2024). and includes locational (i.e., nodal) prices. In this study, 24 weeks spanned across 3 years were analyzed: 1989 (weeks 04, 10, 11, 17, 20, 31, 40, 52), 1995 (weeks 02, 12, 16, 21, 27, 36, 38, 49) and 2009 (weeks 04, 08, 11, 15, 16, 21, 31, 48). For each day in these weeks, nodal prices for every 2-hour interval are available. The file containing the prices for the German-Luxembourgish bidding zone includes 6,090 unique node IDs and 14,170,464 nodal prices. Moreover, for 1,893,024 ($\approx 13\%$) nodal prices, which correspond to 939 distinct nodes, the node ID is “nan”. Consequently, these prices cannot be considered in our experiments, as insufficient data is provided for identifying the nodes to which these prices correspond.
@@
Next, we identify the prices corresponding to the topological nodes. We do this based on the IDs of these nodes and the IDs of the 6,090 nodes included in the file containing LMPs. For 2,200 out of 2,898 topological nodes we could identify their IDs in this file. These nodes belong to 714 substations across Germany. Looking at the different LMP prices computed by ENTSO-E ordered by substation, it turns out that prices for different topological nodes associated with a substation are the same or very close. So, with the 714 substations covered in our study, we could reconstruct most of the German substations, their prices and location across Germany.
@@ 5 Results
In this section, we first analyze the price zone configurations proposed by ACER. Then, we evaluate the optimal price zone configurations computed with K-Means and Spectral Clustering for the years 1989, 1995 and 2009. Considering the metrics defined in Section 3.2, we assess the stability of the resulting clusterings and compare them to those proposed by ACER.
@@ 5.1 Analysis of the price zone configurations proposed by ACER
The configurations proposed by ACER for Germany are the following: two configurations with two price zones (DE2 (k-means) and DE (spectral)), one with three price zones (DE3 (spectral)) and one with four price zones (DE4 (spectral)). Appendix A contains approximations of these configurations given the nodes we consider in our experiments and based on the maps provided by ACER (ACER, 2022a ). Since it is difficult to manually obtain an accurate geographic separation of the zones within the proposed configurations, the configurations from Appendix A are just approximations. In what follows, we analyze these configurations.
@@
Before this, we analyze the nodal prices for 1989, 1995 and 2009 in the configurations proposed by ACER. For each year and week, we compute the cluster price standard deviations and average prices corresponding to configurations DE2 (k-means), DE2 (spectral), DE3 (spectral) and DE4 (spectral), respectively. The purpose of this analysis is to see what benefit multiple price zones have on the reduction of the price standard deviation inside zones. Furthermore, we study whether the clusters are well-defined, i.e., whether, for a given configuration, the cluster average prices are different.
The evolution of cluster price standard deviations across the weeks in 1989, 1995 and 2009 is illustrated in Appendix B.1. We also observe the price standard deviations corresponding to a single price zone or cluster (Figure 5). Clearly, the average cluster price standard deviations decrease as we consider configurations with 2, 3 or 4 clusters. This is in fact an implicit objective of clustering nodes based on prices. However, this reduction is not substantial. For instance, for 2009, the percentage decreases in average price standard deviations compared to a single zone range from 1.92 to 5.69 depending on the configuration. Moreover, the decrease in average cluster price standard deviation is not monotonic with respect to the number of clusters. That is, a larger number of clusters does not necessarily imply smaller average cluster price standard deviations. For instance, for 2009 and configuration DE2 (k-means), the average price standard deviation is 10.85, while DE3 (spectral) leads to an average price standard deviation of around 11.26. An explanation for this behavior is that the average cluster prices are heavily influenced by outliers, which are nodes having very large or low prices compared to the majority of nodes. For smaller clusters, that is, for configurations with a larger number of clusters, the influence of outliers is stronger. Moreover, if we compare the configuration with one price zone to any configuration with multiple zones (clusters), we notice that a reduction in average price standard deviation does not necessarily imply that all cluster standard deviations are lower.
@@
Let us now compare the configurations proposed by ACER to the clusterings computed with our implementation of K-Means and Spectral Clustering. Given the nodal prices for 1989, 1995 and 2009, respectively, we apply both clustering algorithms to identify clusterings (i.e., price zones) for each year. The goal of this comparison is to understand whether the proposed configurations by ACER are stable over time. We say that the ACER configurations are stable if they are similar to the optimal clusterings identified for different years.
As described in Section 4.1, we compute clusterings for two scenarios. The first scenario uses, as clustering features, the time series of nodal prices, while the second scenario additionally considers the location coordinates. Table 1 shows Rand indices measuring the similarity between the clusterings proposed by ACER and the clusterings computed with each algorithm for 1989, 1995 and 2009.
@@
K-Means DE2 (k-means) 0.5 0.5 0.51 0.5 0.5 0.51
DE2 (spectral) 0.5 0.5 0.5 0.5 0.5 0.5
DE3 (spectral) 0.51 0.53 0.52 0.52 0.52 0.52
DE4 (spectral) 0.6 0.5 0.56 0.6 0.58 0.57
Spectral DE2 (k-means) 0.51 0.51 0.51 0.51 0.51 0.51
DE2 (spectral) 0.5 0.5 0.5 0.5 0.5 0.5
DE3 (spectral) 0.52 0.44 0.52 0.52 0.44 0.52
DE4 (spectral) 0.55 0.45 0.57 0.56 0.46 0.56
Table 1: Comparison of alternative price configurations to the price zones proposed by ACER. The table shows Rand indices, which are in $[0,1]$, 1 indicating perfect similarity. The clusterings analyzed in this table were computed with K-Means or Spectral Clustering using time series of nodal prices (columns “Without loc.”) or time series of nodal prices and the location coordinates of nodes (columns “With loc.”).

### v4

# Challenges in Finding Stable Price Zones in European Electricity Markets: Aiming to Square the Circle?

@@ Abstract
The European day-ahead electricity market is split into multiple bidding zones with a uniform price. The increase in renewables leads to a growing number of interventions in the generation of energy sources and increasing redispatch costs. To ensure efficient congestion management, the EU Commission mandated a Bidding Zone Review (BZR) to reevaluate the configuration of European bidding zones. An integral part of this process was a locational marginal pricing study. Based on these prices, alternative bidding zone configurations were proposed. These bidding zones shall be stable and robust over time. For Germany, four configurations were suggested. We analyzed the proposed configurations considering different clustering algorithms and periods based on the publicly released data set in the context of the BZR, and found that the configurations do not reduce the price standard deviations within zones much, and the average prices across zones are similar. Other configurations identified based on clustering the prices lead to lower price variance but they are not geographically coherent. Independent of the clustering features and algorithms used, the resulting clusters are not stable over time. Interestingly, the effect of a split on prices would be low based on an analysis of the BZR data set.
@@ 1 Introduction
The European electricity market is split into multiple bidding zones (BZs) or price zones. Within these zones, market participants can submit bids and offers, which are subsequently matched without considering the capacities of the transmission network. [^1] We will use the terms price zone and bidding zone interchangeably. As shown in Figure 1, BZs are generally defined depending on national borders, such that a country represents a single bidding zone (e.g., France and Poland). However, there are also countries that are divided into several BZs, such as Norway, Sweden and Italy, or BZs that consist of more than one country, like the German-Luxembourgish bidding zone.
Figure 1: European Bidding Zones (adapted from Florence School of Regulation, (2020))
The implementation of European bidding zones began with the First Energy Package in the late 1990s, which aimed to liberalize national electricity markets and set the stage for cross-border trading (Meeus, 2020). This was followed by the Second Energy Package in 2003, which further opened the markets and led to the organization of national electricity grids into bidding zones, often aligned with national borders. In 2006, the Nordic region pioneered market coupling, integrating Norway, Sweden, Finland, and Denmark into a regional electricity market with distinct bidding zones. The Third Energy Package of 2009 formalized the concept of cross-border bidding zones as part of creating a single European electricity market. Most recently, the introduction of Network Codes (Commission, 2013) aimed to harmonize electricity market operations across Europe, enhancing cross-border coordination and the efficiency of bidding zones.
@@
An important characteristic of European BZs is that in the day-ahead market, they implement a uniform pricing mechanism, meaning that a single market-clearing electricity price applies to all participants belonging to a BZ. Europe’s adoption of zonal pricing over the more granular locational marginal pricing (LMP) model seen in markets like PJM reflects a balance between economic efficiency, simplicity, and political practicality (Meeus, 2020, Knörr et al., 2024, Bjorndal & Jornsten, 2001). While LMP offers greater efficiency by reflecting real-time grid conditions at each node, it introduces operational complexity and price volatility. In contrast, the European approach of implementing a uniform price in a whole BZ promotes transparency due to a single price signal in a large region, and it limits the market power that individual parties can exercise. Moreover, this approach enhances liquidity and makes cross-border trading easier in a multi-country system (NEMO Committee, 2024). However, a uniform electricity price also leads to significant negative ramifications. There are no local incentives to adapt to changes in the volatile supply of renewables, there are significant welfare losses due to the uniform price constraint, and large redispatch costs to resolve internal congestion. The latter is because transmission constraints are ignored within a BZ when supply is matched with demand (Energiewende Team, 2018), and, therefore, the market-based allocation is not physically feasible. In addition, uniform prices do not set appropriate investment incentives.
@@
As an example, in the German-Luxembourgish BZ strong winds along the North Sea and Baltic Coast can contribute to elevated power production from wind farms in northern Germany. The energy produced in the north must be transported to the demand centers in the south of the country. However, the transmission lines connecting these regions might become congested and thus electricity cannot be delivered as determined in the day-ahead market. To stabilize the grid, redispatch by the Transmission System Operators (TSOs) is required. TSOs often have to throttle the generation of northern conventional and wind power plants. At the same time, they instruct power plants in the south to increase energy production. Such redispatch actions incur substantial costs. In 2023, the redispatch costs in Germany were 3.1 billion Euros. [^2] https://www.bundesnetzagentur.de/
1.1 The Bidding Zone Review
@@
Based on this ACER methodology (ACER, 2020), ENTSO-E was tasked to calculate locational marginal prices (essentially, a nodal price for each substation in the transmission grid [^4] https://www.tennet.eu/news/bidding-zone-review-tsos-investigate-alternative-bidding-zone-configurations) for the target year 2025, for which generation and demand scenarios were created. These scenarios originate from ENTSO-E’s Data & Models working group and were obtained following the procedure described in Annex 1 of the LMP Study Report (ENTSO-E, 2022). Weather data from the most representative combination of three years in terms of occurrence of residual loads [^5] The residual load is defined as total demand minus renewable generation. was used to create scenarios. This combination was identified considering the years between 1987 and 2016 by an algorithm that used the solar and wind infeed, as well as the hydro inflows and the load as variables. The algorithm selected the combination for which the aggregate dataset of hourly residual loads has the highest average mean and standard deviation relative to all other combinations. Eight representative weeks were identified for each year, with at least one week per season. The grid model, the generation capacities and the load were predicted for 2025 based on several data sources and models which are shown in Figure 2.
Figure 2: Overview of data sources and models used in the LMP Study
ENTSO-E computed bihourly nodal prices for the selected eight weeks in a report published in June 2022 (ENTSO-E, 2022). The prices were obtained based on a linear unit commitment model. As outlined above (ACER, 2022b ), the time series of nodal prices represented the clustering features that were used to partition the set of nodes in the transmission network into a predefined number of clusters or BZs. A BZ configuration or a clustering thus represents the assignment of every node to a specific cluster or price zone. Finally, ACER suggested four alternative bidding zone configurations for Germany ACER, 2022a based on the results from the Locational Marginal Pricing (LMP) Study (ENTSO-E, 2022, 2023). The evaluation of these configurations proposed by ACER in the BZR is ongoing at the time of writing. In this process, the TSOs are evaluating 22 indicators related to market efficiency, network security, and the stability and robustness of the proposed configurations over time (ACER, 2020).
1.2 Contributions
@@
In this paper, we leverage the data from the LMP study provided by ENTSO-E and study how stable and robust the configurations of price zones (or, clusterings) proposed by ACER are with respect to different time frames, different inputs (i.e., nodal prices with or without geographic coordinates), and different clustering algorithms. We define three criteria for analyzing clustering stability and robustness and conduct a thorough empirical evaluation of the suggested configurations using the data published by ENTSO-E. Following ACER’s methodology, we implement K-Means and Spectral Clustering and evaluate the clusterings computed with these methods. We want to understand if the clustering algorithms compute the same clusters when nodal prices of different time periods are considered. How different are clusters resulting from different clustering algorithms? Also, how similar are the clusters of a particular configuration with respect to prices, and how large are the price standard deviations within clusters?
@@
We find that the configurations proposed by ACER are not stable across time and the algorithm used. Moreover, the configurations do not reduce the price standard deviations within zones considerably, and the average prices between zones are similar. If we recompute the clusters and take only prices into account [^6] This means that we do not include constraints in the clustering computation such that we obtain balanced clusters in terms of either the number of nodes or geographical coverage., we can reduce the price standard deviation within clusters further. However, the resulting clusters are not of equal size and are not geographically coherent. Importantly, neither the configurations proposed by ACER nor the clusters that we computed are stable across time. Depending on the time frame taken into account, different clusterings would result.
The data from the LMP study of the BZR can be considered the most comprehensive data set based on up-to-date information from European TSOs. Based on this data set, splitting Germany into 2, 3 or 4 zones does not lead to large price differences between zones and does not reflect long-term price patterns. Our study analyzes the stability and robustness of the alternative configurations with respect to prices, while other factors, such as network security and the impact on energy transition (ACER, 2020), are relevant to evaluate the positive and negative effects of a bidding zone split.
There are different reasons for or against a split of the German price zone. The method adopted in the BZR is based on prices from an LMP study and clustering, and it aims for robust and stable zones. This paper does not make an argument for or against a zonal split. However, it shows that finding geographically large and stable price zones might not be possible. For the question of whether to split the German price zone and how these price zones should be shaped, the clustering of LMPs allows for many answers.
@@ 2 Related Literature
This paper complements existing literature computing bidding zone configurations. In particular, we draw on a unique data set with LMP prices that is provided by ENTSO-E. Bovo et al., (2019) identifies the main network indicators that should be considered for defining suitable bidding zone configurations. LMPs and power transfer distribution factors (PTDFs) represent such indicators. While some papers focus on identifying configurations by clustering PTDFs (e.g., (Kumar et al., 2004) and (Kłos et al., 2014)), most works use LMPs as clustering features (Stoft, 1997). Burstedde, (2012) computes LMPs for the scenario years 2015 and 2020 considering a simplified 72-node representation of the European grid, and evaluates the shapes and sizes of the zones obtained by clustering these LMPs using hierarchical clustering. Breuer et al., (2013) implement a genetic algorithm and apply it to a dataset consisting of 380 kV nodes considering a projected system for 2016 and 2018. Based on six scenarios, Felling & Weber, (2018) compute robust price zone configurations for Central Western Europe by employing hierarchical clustering based on the LMPs of approximately 2,200 nodes and assigning weights to nodes depending on their demand- and supply situation. These configurations are robust in the sense that LMPs from multiple scenarios are integrated into their computation. Considering a network model consisting of around 2,700 nodes, Brouhard et al., (2023) design nine zone delineations for Continental Europe by implementing K-Means and hierarchical clustering. The authors calculate LMPs for multiple, independent single-period optimization problems (i.e., optimal power flow with direct current approximation), and use geographic coordinates and these prices as clustering features.
Sarfati et al., (2015) introduce five indicators to assess the impact of bidding zone configurations in zonally-priced electricity markets. The indicators include commercial exchange and welfare evolution, as well as price convergence and divergence, and were analyzed on the Nordic 32-bus example system. Using a five-node example with a single congested line, Bjorndal & Jornsten, (2001) show that defining zones based on nodal prices is not guaranteed to result in a zonal configuration with maximum welfare. The authors formulate a mixed-integer nonlinear program that, for a given number of zones, computes a zonal configuration with maximum welfare.
@@
In terms of the German-Luxembourgish bidding zone, Zinke, (2023) identifies north-south splits using hierarchical clustering based on the LMPs associated to approximately 500 nodes that are computed within a linear market and grid model for four scenario years. The author considers uncertain factors such as short-term weather patterns and long-term system changes to evaluate the potential decrease in redispatch costs resulting from such splits. Moreover, Ambrosius et al., (2020) consider a scenario for the year 2035 and determine welfare-maximizing price zones and available transfer capacities for a simplified representation of the German market area consisting of 28 nodes. Such zones are identified based on a mixed-integer nonlinear model that incorporates a graph partitioning problem on the first level (Grimm et al., 2019, Kleinert & Schmidt, 2019). The prior literature used different data sources for these analyses.
@@ 3.1 Clustering Algorithms
Following ACER’s approach to identify alternative price zone configurations, we consider two clustering algorithms in this paper. ACER used three clustering approaches (ACER, 2022b ), i.e., constrained K-Means, Spectral Clustering with constrained K-Means and Spectral Clustering with CDIRC, but the implementation details are not publicly available. Therefore, it is impossible to reproduce the exact algorithmic configuration ACER implemented. Instead, we consider the same family of clustering methods to evaluate alternative price zone configurations. In what follows, we briefly describe the clustering algorithms we implement.
3.1.1 Constrained K-Means
@@
3.1.2 Spectral Clustering with Constrained K-Means
@@ 3.2.2 Temporal Stability
While geographic coordinates are static across time, prices are time-sensitive. Therefore, we are interested to see if pricing zones are stable over time. To decide this, we compare two clusterings $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$ that are computed for two different time periods with the same algorithm and considering the same features. For this comparison, we use the Rand index, which measures the similarity between $\mathcal{C}_{1},\mathcal{C}_{2}$ considering all pairs of data points and counting the pairs that are assigned to the same or different clusters in both $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$:
@@ 4 Data
The data we use in our experiments was published as part of the locational pricing study conducted by ENTSO-E [^9] The datasets can be accessed at https://www.entsoe.eu/network_codes/bzr/ (January 2024). and includes locational (i.e., nodal) prices. In this study, 24 weeks spanned across 3 scenario years were analyzed: 1989 (weeks 04, 10, 11, 17, 20, 31, 40, 52), 1995 (weeks 02, 12, 16, 21, 27, 36, 38, 49) and 2009 (weeks 04, 08, 11, 15, 16, 21, 31, 48). For each day in these weeks, nodal prices for every 2-hour interval are available. The file containing the prices for the German-Luxembourgish bidding zone includes 6,090 unique node IDs and 14,170,464 nodal prices. Moreover, for 1,893,024 ($\approx 13\%$) nodal prices, which correspond to 939 distinct nodes, the node ID is “nan”. Consequently, these prices cannot be considered in our experiments, as insufficient data is provided for identifying the nodes to which these prices correspond.
@@
Next, we identify the prices corresponding to the topological nodes. We do this based on the IDs of these nodes and the IDs of the 6,090 nodes included in the file containing LMPs. For 2,200 out of 2,898 topological nodes we could identify their IDs in this file. These nodes belong to 714 substations across Germany. Looking at the different LMP prices computed by ENTSO-E ordered by substation, it turns out that prices for different topological nodes associated with a substation are the same or very close. So, with the 714 substations covered in our study, we could reconstruct most of the German substations, their prices and locations across Germany.
@@ 5 Results
In this section, we first analyze the price zone configurations proposed by ACER. Then, we evaluate the optimal price zone configurations computed with our implementation of K-Means and Spectral Clustering for the scenario years 1989, 1995 and 2009. Considering the metrics defined in Section 3.2, we assess the stability of the resulting clusterings and compare them to those proposed by ACER. In what follows, whenever we mention K-Means and Spectral Clustering as clustering methods, we refer to our implementation described in Section 3.1.
@@ 5.1 Analysis of the price zone configurations proposed by ACER
The configurations proposed by ACER for Germany are specified in Annex I of ACER’s Decision on the alternative bidding zone configurations to be considered in the bidding zone review process (ACER, 2022a ). For brevity reasons, we refer to the two-zone configurations computed using K-Means and Spectral Clustering as DE2 (k-means) and DE2 (spectral), respectively. Similarly, the configurations with three and four zones are denoted as DE3 (spectral) and DE4 (spectral). Appendix A contains approximations of these configurations given the nodes we consider in our experiments and based on the maps provided by ACER (ACER, 2022a ). Since it is difficult to manually obtain an accurate geographic separation of the zones within the proposed configurations, the configurations from Appendix A are just approximations. In what follows, we analyze these configurations.
@@
Before this, we analyze the nodal prices for scenario years 1989, 1995 and 2009 in the configurations proposed by ACER. For each year and week, we compute the cluster price standard deviations and average prices corresponding to configurations DE2 (k-means), DE2 (spectral), DE3 (spectral) and DE4 (spectral), respectively. The purpose of this analysis is to see what benefit multiple price zones have on the reduction of the price standard deviation inside zones. Furthermore, we study whether the clusters are well-defined, i.e., whether, for a given configuration, the cluster average prices are different.
The evolution of cluster price standard deviations across the weeks in the scenario years 1989, 1995 and 2009 is illustrated in Appendix B.1. We also observe the price standard deviations corresponding to a single price zone or cluster (Figure 8). Clearly, the average cluster price standard deviations decrease as we consider configurations with 2, 3 or 4 clusters. This is an implicit objective of clustering nodes based on prices. However, this reduction is not substantial and varies between 3% to 9% depending on the configuration. For instance, in 2009, the percentage decreases in average price standard deviations compared to a single zone range from 1.92 to 5.69. Moreover, the decrease in average cluster price standard deviation is not monotonic with respect to the number of clusters. That is, a larger number of clusters does not necessarily imply smaller average cluster price standard deviations. For example, for 2009 and configuration DE2 (k-means), the average price standard deviation is 10.85, while DE3 (spectral) leads to an average price standard deviation of around 11.26. An explanation for this behavior is that the average cluster prices are heavily influenced by outliers, which are nodes having very large or low prices compared to the majority of nodes. For smaller clusters, that is, for configurations with a larger number of clusters, the influence of outliers is stronger. Moreover, if we compare the configuration with one price zone to any configuration with multiple zones (clusters), we notice that a reduction in average price standard deviation does not necessarily imply that all cluster standard deviations are lower.
@@
Let us now compare the configurations proposed by ACER to the clusterings computed with our implementation of K-Means and Spectral Clustering. Given the nodal prices for the scenario years 1989, 1995 and 2009, respectively, we apply both clustering algorithms to identify clusterings (i.e., price zones) for each year. The goal of this comparison is to understand whether the proposed configurations by ACER are stable over time. We say that the ACER configurations are stable if they are similar to the optimal clusterings identified for different years.
As described in Section 4.1, we compute clusterings for two scenarios. The first scenario uses, as clustering features, the time series of nodal prices, while the second scenario additionally considers the location coordinates. Table 1 shows Rand indices measuring the similarity between the clusterings proposed by ACER and the clusterings we computed with each algorithm for the scenario years 1989, 1995 and 2009. For example, the first Rand index (0.5, column 3) indicates the similarity between (1) the solution computed with our implementation of K-Means using the time series of nodal prices corresponding to the scenario year 1989 and (2) configuration DE2 (k-means) proposed by ACER.
@@
K-Means ACER DE2 (k-means) 0.5 0.5 0.51 0.5 0.5 0.51
ACER DE2 (spectral) 0.5 0.5 0.5 0.5 0.5 0.5
ACER DE3 (spectral) 0.51 0.53 0.52 0.52 0.52 0.52
ACER DE4 (spectral) 0.6 0.5 0.56 0.6 0.58 0.57
Spectral ACER DE2 (k-means) 0.51 0.51 0.51 0.51 0.51 0.51
ACER DE2 (spectral) 0.5 0.5 0.5 0.5 0.5 0.5
ACER DE3 (spectral) 0.52 0.44 0.52 0.52 0.44 0.52
ACER DE4 (spectral) 0.55 0.45 0.57 0.56 0.46 0.56
Table 1: Comparison between the alternative price configurations computed with our implementation of K-Means and Spectral Clustering and the price zones proposed by ACER. The table shows Rand indices, which are in $[0,1]$, 1 indicating perfect similarity. We computed clusterings using K-Means and Spectral Clustering using: (1) time series of nodal prices (columns “Without loc.”) and (2) time series of nodal prices and the location coordinates of nodes (column “With loc.”).